# MotionLab: Unified Human Motion Generation and Editing via the Motion-Condition-Motion Paradigm

### v3

@@ Abstract.
Human motion generation and editing are key components of computer graphics and vision. However, current approaches in this field tend to offer isolated solutions tailored to specific tasks, which can be inefficient and impractical for real-world applications. While some efforts have aimed to unify motion-related tasks, these methods simply use different modalities as conditions to guide motion generation. Consequently, they lack editing capabilities, fine-grained control, and fail to facilitate knowledge sharing across tasks. To address these limitations and provide a versatile, unified framework capable of handling both human motion generation and editing, we introduce a novel paradigm: Motion-Condition-Motion, which enables the unified formulation of diverse tasks with three concepts: source motion, condition, and target motion. Based on this paradigm, we propose a unified framework, MotionLab, which incorporates rectified flows to learn the mapping from source motion to target motion, guided by the specified conditions. In MotionLab, we introduce the 1) MotionFlow Transformer to enhance conditional generation and editing without task-specific modules; 2) Aligned Rotational Position Encoding to guarantee the time synchronization between source motion and target motion; 3) Task Specified Instruction Modulation; and 4) Motion Curriculum Learning for effective multi-task learning and knowledge sharing across tasks. Notably, our MotionLab demonstrates promising generalization capabilities and inference efficiency across multiple benchmarks for human motion. Our code and additional video results are available at: Project Website.
@@ 1. Introduction
Human motion is a crucial component of computer graphics and vision, with applications spanning game development, film production, and virtual reality (Guo et al., 2022a; Tevet et al., 2023). With the advancements of generative diffusion models (Ho et al., 2020; Song et al., 2020; Dhariwal and Nichol, 2021), human motion generation has garnered considerable attention, aiming at generating human motion aligned with the input conditions, such as text (Tevet et al., 2023; Zhang et al., 2022; Tevet et al., 2022) and trajectory (i.e., joints’ coordinates) (Xie et al., 2023; Fujiwara et al., 2025; Sun et al., 2024; Dai et al., 2025; Athanasiou et al., 2023, 2022; Zhang et al., 2023a; Song et al., 2023; Sampieri et al., 2024). On the other hand, in order to modify the motion assets in the industry, significant efforts have been dedicated to motion editing, which intended to modify the properties of prepared motion like motion style transfer (Zhong et al., 2025; Song et al., 2024; Aberman et al., 2020; Guo et al., 2024b; Jang et al., 2022).
@@
As illustrated in Table 1, existing research in this domain predominantly focuses on isolated solutions tailored to specific tasks. Consequently, users are compelled to train multiple distinct models for generating and editing human motion, which is both inefficient and impractical. Although some studies (Shrestha et al., 2025; Zhou and Wang, 2023; Zhou et al., 2023; Zhang et al., 2025; Yang et al., 2024; Luo et al., 2024; Fan et al., 2024) have endeavored to unify motion-related tasks, they merely treat various modalities as conditions for motion generation and naively integrate them into a single model. Such approaches not only constrain their editing capabilities but also fail to provide fine-grained trajectory control. Furthermore, these solutions fail to leverage the intrinsic connections between motion generation and editing, thereby hindering knowledge sharing across tasks. Instead, a unified framework of generation and editing offers the potential to expand data volume by incorporating datasets from different tasks, thereby enhancing model performance on data-scarce tasks. Motivated by this benefit, as well as the success of large language models (LLMs)(Achiam et al., 2023; Dubey et al., 2024) in unifying diverse tasks to significantly enhance capability and generalization, we pose the following question: Can human motion generation and editing be effectively unified within a single framework?
@@
To achieve this, it is essential to design an elegant and scalable paradigm. Drawing inspiration from the next-token prediction paradigm (Brown et al., 2020; Achiam et al., 2023), which has revolutionized the NLP field, we propose a novel paradigm: Motion-Condition-Motion. This paradigm is built upon three concepts – source motion, condition, and target motion. Similar to how next-token prediction anticipates the subsequent word based on context, the Motion-Condition-Motion paradigm predicts the target motion based on the source motion and specified conditions. For any human motion generation task, the source motion can be treated as none, and the target motion must align with the provided conditions. For any human motion editing task, the target motion is derived from the source motion based on the conditions. By unifying these tasks within this elegant and scalable paradigm, this framework can be seamlessly extended to various human motion tasks and scaled across diverse datasets. Given that human motions are inherently tied to their semantics, trajectories, and styles in practical applications, we aim to unify several key tasks under this framework. These tasks include text-based motion generation and editing (Tevet et al., 2023; Zhang et al., 2022; Tevet et al., 2022; Athanasiou et al., 2024), trajectory-based motion generation and editing (Xie et al., 2023; Dai et al., 2025), motion in-between (Cohan et al., 2024) and motion style transfer (Song et al., 2024; Zhong et al., 2025), as illustrated in Figure 1.
@@
Despite the proposed paradigm, significant challenges remain in balancing versatility, performance, and efficiency: 1) Unifying various tasks inevitably introduces additional modalities. However, adopting multiple cross-attention mechanisms for each task, as employed in generation-unified frameworks (Fan et al., 2024; Zhang et al., 2025), is both inefficient and suboptimal. This cross-attention approach risks neglecting the rich, deep features of these modalities and fails to establish direct interactions between the modalities and the target motion. 2) More sampling time is required for certain tasks (e.g., trajectory-based motion generation and motion in-between (Xie et al., 2023; Zhong et al., 2025)), as existing methods in these areas involve task-specific posterior guidance (Chung et al., 2022) during inference to improve conditional guidance. 3) Time asynchrony between the source motion and target motion may arise due to the limited scale of the paired editing dataset and the use of implicit positional encoding (Chen et al., 2023; Xie et al., 2023; Athanasiou et al., 2024; Cohan et al., 2024). 4) Most importantly, naively integrating various motion generation and editing tasks into a single framework can lead to task conflicts and catastrophic forgetting, impairing the framework’s overall performance.
@@
To address these challenges, we propose a powerful generative framework, named MotionLab, built upon our designed MotionFlow Transformer (MFT) as shown in Figure 3. Inspired by MM-DiT (Esser et al., 2024), our MFT also leverages rectified flows (Liu et al., 2022; Lipman et al., 2022), but we utilize them to map source motion to target motion based on specified conditions. Furthermore, unlike MM-DiT, which focuses exclusively on text and images, our MFT incorporates multiple modalities: source motion, target motion, text, trajectory, and style. In MFT, each modality is equipped with its own conditioning path for deep feature and fully interacts with the others through joint attention. This design enables MFT to enhance conditional generation and editing without requiring task-specific modules or posterior guidance for certain tasks. To ensure temporal synchronization between source motion and target motion, we incorporate an Aligned ROtational Position Encoding (Aligned ROPE) into MFT, explicitly aligning tokens in corresponding frames between the source and target motion. Additionally, to harmoniously integrate diverse tasks, we propose Task Instruction Modulation, which distinguishes tasks by introducing an additional embedding representing the task into the MFT.
To facilitate effective multi-task training, we have devised a curriculum-inspired training strategy, termed Motion Curriculum Learning, which accounts for the varying difficulty levels of different tasks. Intuitively, tasks involving fewer modalities or clearer condition information are simpler; for example, the reconstruction of source motion is easier than text-based motion generation. Adhering to the easy-to-hard training principle, we initially introduce the masked source motion reconstruction to extract discriminative motion features. Consequently, the training process in MotionLab is divided into two sequential stages: masked pre-training and supervised fine-tuning. During masked pre-training, the model reconstructs masked source motion to establish a foundational understanding of motion features. In the supervised fine-tuning stage, the model progressively learns new tasks in an easy-to-hard sequence. Our Motion Curriculum Learning not only fosters harmonious knowledge sharing across tasks but also supports Classifier-Free Guidance (CFG) (Ho and Salimans, 2022) during inference, resulting in improved performance across a variety of tasks.
The main contributions of this work are as follows: 1) We introduce the Motion-Condition-Motion paradigm, which reformulates human motion generation and editing tasks using three core concepts, enabling a unified formulation. 2) We propose the MotionLab framework within the Motion-Condition-Motion paradigm, featuring MotionFlow Transformer to enhance the conditional generation and editing, Aligned ROPE to ensure the synchronization between source motion and target motion, and Task Instruction Modulation for effective multi-task learning. 3) We develop the Motion Curriculum Learning strategy, which enables hierarchical training from pre-training to sequential fine-tuning, effectively sharing the knowledge between various tasks. 4) We validate MotionLab on multiple benchmarks, demonstrating its superior versatility, performance, and efficiency over baselines.
@@ 2. Related Works
Motion Generation and Editing. Motion generation can be classified based on input conditions. Among these, text-based motion generation is one of the most compelling areas (Zhang et al., 2022; Tevet et al., 2022, 2023; Guo et al., 2022a; Chen et al., 2023; Guo et al., 2024a, 2022c; Wang et al., 2024b; Jiang et al., 2023; Kim et al., 2023; Lin et al., 2023; Lu et al., 2023; Plappert et al., 2016; Zhang et al., 2023b; Guo et al., 2020; Petrovich et al., 2021), as it trains models to comprehend the semantics of text and generate corresponding pose sequences. To address the fine-grained requirements of practical applications, trajectory-based motion generation has been proposed (Karunratanakul et al., 2023; Shafir et al., 2023; Xie et al., 2023; Dai et al., 2025; Zhang et al., 2023a), where specific motion properties, such as joints reaching designated positions at specified times, are defined. Additionally, motion in-between (Tevet et al., 2023; Jiang et al., 2023; Cohan et al., 2024; Qin et al., 2022; Pinyoanuntapong et al., 2024) focuses on generating complete motion sequences given key poses at keyframes. To enable in-place editing of human motion, MotionFix (Athanasiou et al., 2024) introduces text-based motion editing using paired source and target motions. We extend this approach to trajectory-based motion editing by substituting text with joint trajectories. Meanwhile, style plays a crucial role in human motion, leading to motion style-transfer (Jang et al., 2022; Aberman et al., 2020; Zhong et al., 2025; Song et al., 2024). However, the aforementioned methods concentrate solely on specific tasks, rendering them impractical for real-world applications. Moreover, they overlook the intrinsic connections across different human motion tasks and fail to facilitate knowledge sharing among these tasks. In contrast, our unified framework enhances performance on data-scarce editing tasks through multi-task learning.
@@
Unified frameworks for human motion. There are also some efforts in existing methods that try to unify tasks related to human motion. One line of work (Jiang et al., 2023, 2025; Zhou et al., 2024; Li et al., 2024; Wang et al., 2024b; Ling et al., 2024; Wu et al., 2024; Luo et al., 2024) focuses on motion understanding, such as motion captioning or describing human motion in images and videos. Yet, these approaches often rely on GPT-like structures, which requires a large amount of training resources and GPU memory. In addition, they fail to provide fine-grained control (e.g., trajectory-based generation and editing) over motion, which is crucial in practical applications. Another line of effort (Shrestha et al., 2025; Zhou and Wang, 2023; Zhou et al., 2023; Zhang et al., 2025; Yang et al., 2024; Luo et al., 2024; Fan et al., 2024) highlights generating motion based on more modalities, such as music and speech. However, these approaches just only integrate more modalities into one model and cannot flexibly edit motion, which can cause them suffering from the multi-task learning and limit their scope of use. The closest to our work is MotionCLR (Chen et al., 2024), which does not support fine-grained control (e.g., trajectory-based generation and editing) and requires cumbersome manual adjustments to the attention mechanism.
@@ 3. PRELIMINARY: Rectified Flows
Flow matching methods (Liu et al., 2022; Lipman et al., 2022; Esser et al., 2024; Ma et al., 2024; Fei et al., 2024; Polyak et al., 2024) have recently received significant attention due to their generalizability and efficiency. As shown in Figure 2, these methods directly regress the transport vector field between noise distribution $p_{1}$ and data distribution $p_{0}$ with the straightest possible trajectories and sample by the corresponding ordinary differential equation (ODE) (Wang et al., 2024a). Among these methods, rectified flows (Liu et al., 2022; Lipman et al., 2022) aim to learn a trajectory from noise $x_{0}$ to data $x_{1}$, which can be formulated as $x_{t}=\varphi(x_{0},x_{1},t)$, and the velocity field $v_{t}$ of the trajectory $x_{t}$ can be defined by:
@@ 4. Motion-Condition-Motion
Remarks. In particular, trajectory-based motion generation and motion in-between are highly similar, as they both aim to ensure that specific joints reach designated positions at specific times. The primary distinction between them is that the former is sparse in space (i.e., joints) but dense in time, whereas the latter is dense in space (i.e., joints) but sparse in time. To efficiently share the parameters and learned representations between the two tasks, we unify their conditions into a single condition. Meanwhile, masked reconstruction is also similar to these two tasks. However, while these two tasks only include the coordinates of joints, the source motion also encompasses the velocity and angular velocity of joints. Therefore, they represent different modalities, and masked reconstruction constitutes a distinct task.
@@ 5.1. MotionFlow Transformer
As shown in the Figure 3 (b), MotionFlow Transformer contains three key components: Joint Attention to interact tokens from different modalities; Condition Path for distinguishing tokens from different modalities and extracting their representations, and Aligned ROPE for position encoding of modalities with time information.
@@
Condition Path. While the joint attention is able to interact tokens from different modalities, there is still need to differentiate different tokens. In addition to the QKV projection and FeedForward Network (FFN) in the attention mechanism, as used in MM-DiT, our MFT incorporates the adaptive Layer Normalization (adaLN) and a modulation mechanism (Peebles and Xie, 2023) for each modality, enhancing conditional generation and editing capabilities.

### v4

@@ Abstract.
Human motion generation and editing are key components of computer vision. However, current approaches in this field tend to offer isolated solutions tailored to specific tasks, which can be inefficient and impractical for real-world applications. While some efforts have aimed to unify motion-related tasks, these methods simply use different modalities as conditions to guide motion generation. Consequently, they lack editing capabilities, fine-grained control, and fail to facilitate knowledge sharing across tasks. To address these limitations and provide a versatile, unified framework capable of handling both human motion generation and editing, we introduce a novel paradigm: Motion-Condition-Motion, which enables the unified formulation of diverse tasks with three concepts: source motion, condition, and target motion. Based on this paradigm, we propose a unified framework, MotionLab, which incorporates rectified flows to learn the mapping from source motion to target motion, guided by the specified conditions. In MotionLab, we introduce the 1) MotionFlow Transformer to enhance conditional generation and editing without task-specific modules; 2) Aligned Rotational Position Encoding to guarantee the time synchronization between source motion and target motion; 3) Task Specified Instruction Modulation; and 4) Motion Curriculum Learning for effective multi-task learning and knowledge sharing across tasks. Notably, our MotionLab demonstrates promising generalization capabilities and inference efficiency across multiple benchmarks for human motion. Our code and additional video results are available at: Project Website.
@@ 1. Introduction
Human motion is a crucial component of computer vision, with applications spanning game development, film production, and virtual reality (Guo et al., 2022a; Tevet et al., 2023). With the advancements of generative diffusion models (Ho et al., 2020; Song et al., 2020; Dhariwal and Nichol, 2021), human motion generation has garnered considerable attention, aiming at generating human motion aligned with the input conditions, such as text (Tevet et al., 2023; Zhang et al., 2022; Tevet et al., 2022) and trajectory (i.e., joints’ coordinates) (Xie et al., 2023; Fujiwara et al., 2025; Sun et al., 2024; Dai et al., 2025; Athanasiou et al., 2023, 2022; Zhang et al., 2023a; Song et al., 2023; Sampieri et al., 2024). On the other hand, in order to modify the motion assets in the industry, significant efforts have been dedicated to motion editing, which intended to modify the properties of prepared motion like motion style transfer (Zhong et al., 2025; Song et al., 2024; Aberman et al., 2020; Guo et al., 2024b; Jang et al., 2022).
@@
As summarized in Table 1, current research in this domain mainly develops isolated, task-specific solutions, forcing practitioners to train multiple models for human motion generation and editing—a strategy that is inefficient and impractical. Although several studies (Shrestha et al., 2025; Zhou and Wang, 2023; Zhou et al., 2023; Zhang et al., 2025a; Yang et al., 2024; Luo et al., 2024; Fan et al., 2024) have attempted to unify motion-related tasks, they merely consider different modalities as generation conditions, leading to limited editing capabilities and insufficient fine-grained trajectory control. Moreover, these approaches overlook the intrinsic links between motion generation and editing, thereby hindering potential knowledge sharing. In contrast, a well-designed unified framework can exploit the large volumes of multi-task data to potentially surpass specialist models through effective cross-task representation learning. Motivated by this prospect and inspired by the success of large language models in unifying diverse tasks (Achiam et al., 2023; Dubey et al., 2024), we pose the following question: Can human motion generation and editing be effectively unified within a single framework?
@@
In response to this question, we aim to design an elegant and scalable paradigm. Drawing inspiration from the next-token prediction paradigm (Brown et al., 2020; Achiam et al., 2023), which has revolutionized the NLP field, we propose a novel paradigm: Motion-Condition-Motion. This paradigm is built upon three concepts – source motion, condition, and target motion. Similar to how next-token prediction anticipates the subsequent word based on context, the Motion-Condition-Motion paradigm predicts the target motion based on the source motion and specified conditions. For any human motion generation task, the source motion can be treated as none, and the target motion must align with the provided conditions. For any human motion editing task, the target motion is derived from the source motion based on the conditions. By unifying these tasks within this elegant and scalable paradigm, this framework can be seamlessly extended to various human motion tasks and scaled across diverse datasets. Given that human motions are inherently tied to their semantics, trajectories, and styles in practical applications, we aim to unify several key tasks under this framework. These tasks include text-based motion generation and editing (Tevet et al., 2023; Zhang et al., 2022; Tevet et al., 2022; Athanasiou et al., 2024; Goel et al., 2024), trajectory-based motion generation and editing (Xie et al., 2023; Dai et al., 2025), motion in-between (Cohan et al., 2024; Harvey et al., 2020) and motion style transfer (Song et al., 2024; Zhong et al., 2025), as illustrated in Figure 1.
@@
Despite the proposed paradigm, several significant challenges remain in balancing versatility, performance, and efficiency: 1) Unifying various tasks inevitably introduces additional modalities, while each modality may involve multiple tasks. A naive solution, like adopting multiple cross-attention mechanisms for each task in generation-unified frameworks (Fan et al., 2024; Zhang et al., 2025a), is suboptimal. 2) More sampling time is required for certain tasks (e.g., trajectory-based motion generation and motion in-between (Xie et al., 2023; Zhong et al., 2025)), as existing methods in these areas involve task-specific posterior guidance (Chung et al., 2022) during inference to improve conditional guidance. 3) Time asynchrony between the source motion and target motion may arise due to the limited scale of the paired editing dataset and the use of implicit positional encoding (Chen et al., 2023; Xie et al., 2023; Athanasiou et al., 2024; Cohan et al., 2024). 4) Most importantly, naively integrating various motion generation and editing tasks into a single framework could lead to task conflicts and catastrophic forgetting, impairing the framework’s overall performance.
@@
To address these challenges, we propose a novel generative framework, named MotionLab, built upon our designed MotionFlow Transformer (MFT) as shown in Figure 3. Inspired by MM-DiT (Esser et al., 2024), our MFT also leverages rectified flows (Liu et al., 2022; Lipman et al., 2022), but we utilize them to map source motion to target motion based on specified conditions. Unlike MM-DiT that focuses exclusively on text and images, our MFT incorporates multiple modalities: source motion, target motion, text, trajectory, and style. In MFT, each modality is allocated a dedicated modality path and fully interacts with the others through joint attention, which enables MFT to enhance conditional generation and editing without requiring task-specific modules or posterior guidance for certain tasks. To ensure temporal synchronization between source motion and target motion, we incorporate an Aligned ROtational Position Encoding (Aligned ROPE) into MFT, explicitly aligning tokens in corresponding frames between the source and target motion. Additionally, to adapt one modality to different tasks, we propose Task Instruction Modulation, which distinguishes different tasks for each modality by introducing an additional task embedding into the MFT. To harmoniously integrate diverse tasks, we propose a curriculum-inspired training strategy termed Motion Curriculum Learning based on the easy-to-hard training principle. Intuitively, tasks involving fewer modalities or more explicit conditional information (e.g., source motion) present lower complexity and are therefore prioritized in the learning sequence.
In this paper, motion-related tasks are decomposed into combinations of modalities through the Motion-Condition-Motion paradigm. These modalities are subsequently represented via each modality paths within the MFT, learning cross-modal interactions through the joint attention, while adapting individual modalities to different tasks through Task Instruction Modulation. By implementing a curriculum learning from single to multiple, from simple (e.g., source motion and trajectory) to complex (e.g., text and style) modalities, the spatial knowledge inherent in 3D representations can be effectively transferred to more abstract modalities since the modalities of the former can represent the latter. Through these designs, we validate MotionLab on multiple benchmarks, demonstrating superior versatility, performance, and efficiency compared to baselines across various human motion generation and editing tasks.
@@ 2. Related Works
Motion Generation and Editing. Motion generation can be classified based on input conditions. Among these, text-based motion generation is one of the most compelling areas (Zhang et al., 2022; Tevet et al., 2022, 2023; Guo et al., 2022a; Chen et al., 2023; Guo et al., 2024a, 2022b; Wang et al., 2024b; Jiang et al., 2023; Kim et al., 2023; Lin et al., 2023; Lu et al., 2023; Plappert et al., 2016; Zhang et al., 2023b; Guo et al., 2020; Petrovich et al., 2021), as it trains models to comprehend the semantics of text and generate corresponding pose sequences. To address the fine-grained requirements of practical applications, trajectory-based motion generation has been proposed (Karunratanakul et al., 2023; Shafir et al., 2023; Xie et al., 2023; Dai et al., 2025; Zhang et al., 2023a), where specific motion properties, such as joints reaching designated positions at specified times, are defined. Additionally, motion in-between (Tevet et al., 2023; Jiang et al., 2023; Cohan et al., 2024; Qin et al., 2022; Pinyoanuntapong et al., 2024) focuses on generating complete motion sequences given key poses at keyframes. To enable in-place editing of human motion (Goel et al., 2024; Athanasiou et al., 2024), MotionFix (Athanasiou et al., 2024) introduces text-based motion editing using paired source and target motions. We extend this approach to trajectory-based motion editing by substituting text with joint trajectories. Meanwhile, style plays a crucial role in human motion, leading to motion style-transfer (Jang et al., 2022; Aberman et al., 2020; Zhong et al., 2025; Song et al., 2024). However, the aforementioned methods concentrate solely on specific tasks, rendering them impractical for real-world applications. Moreover, they overlook the intrinsic connections across different human motion tasks and fail to facilitate knowledge sharing among these tasks. In contrast, our unified framework enhances performance on data-scarce editing tasks through multi-task learning.
@@
Unified frameworks for human motion. There are also some efforts in existing methods that try to unify tasks related to human motion. One line of work (Jiang et al., 2023, 2025; Zhou et al., 2024; Li et al., 2024; Wang et al., 2024b; Ling et al., 2024; Wu et al., 2024; Luo et al., 2024; Athanasiou et al., 2024) focuses on motion understanding, such as motion captioning or describing human motion in images and videos. Yet, these approaches often rely on GPT-like structures, which requires a large amount of training resources and GPU memory. In addition, they fail to provide fine-grained control (e.g., trajectory-based generation and editing) over motion, which is crucial in practical applications. Another line of effort (Shrestha et al., 2025; Zhou and Wang, 2023; Zhou et al., 2023; Zhang et al., 2025a; Yang et al., 2024; Luo et al., 2024; Fan et al., 2024; Alexanderson et al., 2023) highlights generating motion based on more modalities, such as music and speech. However, these approaches just only integrate more modalities into one model and cannot flexibly edit motion, which can cause them suffering from the multi-task learning and limit their scope of use. The closest to our work are FLAME (Kim et al., 2023) and MotionCLR (Chen et al., 2024). However, FLAME does not support style transfer and precise text-based editing like “move faster”, and MotionCLR does not support trajectory-based generation and editing, requiring cumbersome manual adjustments to the attention.
@@ 3. PRELIMINARY: Rectified Flows
Flow-based methods (Liu et al., 2022; Lipman et al., 2022; Esser et al., 2024; Ma et al., 2024; Fei et al., 2024; Polyak et al., 2024) have recently received significant attention due to their generalizability and efficiency compared to diffusion models. Specifically, these methods directly regress the transport vector field between noise distribution $p_{1}$ and data distribution $p_{0}$ with the straightest possible trajectories and sample by the corresponding ordinary differential equation (ODE) (Wang et al., 2024a). Among these methods, rectified flows (Liu et al., 2022; Lipman et al., 2022) aim to learn a trajectory from noise $x_{0}$ to data $x_{1}$, which can be formulated as $x_{t}=\varphi(x_{0},x_{1},t)$, and the velocity field $v_{t}$ of the trajectory $x_{t}$ can be defined by:
@@ 4. Motion-Condition-Motion
Remarks. In particular, trajectory-based motion generation and motion in-between are highly similar, as they both aim to ensure that specific joints reach designated positions at specific times. Their primary difference is that the former is sparse in space (i.e., joints) but dense in time, whereas the latter is dense in space (i.e., joints) but sparse in time. To efficiently share the parameters and learned representations between the two tasks, we unify their conditions into a single condition. Meanwhile, masked reconstruction is also similar to these two tasks. However, while these two tasks only include the coordinates of joints, the source motion also encompasses the velocity and angular velocity of joints. Therefore, they represent different modalities, and masked reconstruction constitutes a distinct task.
@@ 5.1. MotionFlow Transformer
As shown in the Figure 3 (b), MotionFlow Transformer contains three key components: Joint Attention to interact tokens from different modalities; Modality Path for distinguishing tokens from different modalities and extracting their representations, and Aligned ROPE for position encoding of modalities with time information.
@@
Modality Path. While the joint attention is able to interact tokens from different modalities, there is still need to differentiate different tokens. In addition to the QKV projection and FeedForward Network (FFN) in the attention mechanism, as used in MM-DiT, our MFT incorporates the adaptive Layer Normalization (adaLN) and a modulation mechanism (Peebles and Xie, 2023) for each modality, enhancing conditional generation and editing capabilities.